# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 4.3.2 Speech Recognition
Speech, as the primary mode of human communication, has captivated researchers for over five decades, especially since the inception of artificial intelligence [198]. From the earliest endeavors to understand and replicate the complexities of human speech, to contemporary advancements leveraging cutting-edge technologies, the quest for accurate and efficient speech recognition systems has been relentless. In recent years, the emergence of deep learning techniques has revolutionized the speech recognition field. Deep learning has demonstrated unparalleled success in processing and extracting intricate patterns from vast amount of data. When applied to the realm of speech recognition, deep learning have surpassed traditional approaches by learning intricate features directly from raw audio signals, circumventing the need for handcrafted features and complex preprocessing pipelines. This paradigm shift has significantly advanced the state-of-the-art in speech recognition, enabling systems to achieve unprecedented levels of accuracy and robustness across various languages, accents, and environmental conditions. Generally, deep learning has been extended to other essential applications of speech recognition, such as speaker identification [199, 200], emotion recognition [201], language identification [202], accent recognition [203], age recognition [204] and gender recognition [205], among many others.
@@
In recent times, CNN and RNNs have been leveraged for automatic speech recognition in order to consider a longer or variable temporal window for context information extraction [209]. Generally, CNNs are well-suited for capturing local patterns and hierarchical features in data, making them effective for modelling acoustic features in speech. By directly learning features from raw speech signals, CNNs bypassed the need for handcrafted features used in traditional GMM-HMM systems. Additionally, CNNs can capture long-range dependencies in the data, which is crucial for understanding the context of speech. Likewise, the RNNs are suitable choice for exploring extended temporal context information in one processing level for feature extraction and modelling.
@@
Based on this, several researchers have proposed the use of both CNN and variants of RNNs for automatic speech recognition and for other speech related tasks. For instance, [210], used CNN to classify speech emotions and benchmarked on a dataset consisting of seven classes (anger, disgust, fear, happiness, neutral, sadness and surprise). However, CNN lack the ability to model temporal dependencies explicitly. In speech recognition, understanding the temporal context of speech is essential for accurate transcription. Also, speech signals are inherently sequential, and information from previous time steps is crucial for understanding the current speech segment. CNNs, by design, do not inherently capture this sequential nature. For this reason, variants of RNNs have been leveraged to collect extended contexts in speeches. This is because RNNs are designed to model sequential data by maintaining hidden states that capture information from previous time steps. This allows them to capture temporal dependencies effectively, making them well-suited for ASR tasks. In [211], the authors evaluated the performance of RNN, LSTM, and GRU on a popular benchmark speech dataset (ED-LIUM). The results showed that LSTM achieved the best word error rate while the GRU optimization was faster and achieved word error rate close to that of LSTM.
@@
However, RNN architectures process input sequences sequentially, which limits their ability to capture global context information effectively. As a result, they may struggle to understand the entire context of a spoken utterance, leading to lower transcription accuracy, particularly in tasks requiring understanding beyond local dependencies. Also, most CNN and RNN automatic speech recognition systems comprise of separate acoustic, pronunciation, and language modelling components that are trained independently. Usually, the acoustic model bootstraps from an existing model that is used for alignment in order to train it to recognise context dependent (CD) states or phonemes. The pronunciation model, curated by expert linguists, maps the sequences of phonemes produced by the acoustic model into word sequences. For this reason, Sequence-to-Sequence (Seq2Seq) models are being proposed in automatic speech recognition to train the acoustic, pronunciation, and language modelling components jointly in a single system [212]. Seq2Seq models in automatic speech recognition are a class of models that aim to directly transcribe an input sequence of acoustic features such as speech spectrograms or Mel-frequency cepstral coefficients into a sequence of characters or words representing the recognized speech. There have been a variety of sequence-to-sequence models explored in the literature, including Recurrent Neural Network Transducer (RNN-T) [213], Listen, Attend and Spell (LAS) [214], Neural Transducer [215], Monotonic Alignments [216] and Recurrent Neural Aligner (RNA) [217].
@@
In [220], the authors explored various structural and optimization enhancements to their LAS Sequence to Sequence model, resulting in significant performance improvements. They introduce several structural enhancements, including the utilization of word piece models instead of graphemes and the incorporation of a multi-head attention architecture, which outperforms the commonly used single-head attention mechanism. Additionally, they investigate optimization techniques such as synchronous training, scheduled sampling, label smoothing, and minimum word error rate optimization, all of which demonstrate improvements in accuracy. The authors present experimental results utilizing a unidirectional LSTM encoder for streaming recognition. On a 12,500-hour voice search task, they observe a decrease in Word Error Rate (WER) from 9.2% to 5.6% with the proposed changes, while the best-performing conventional system achieves a WER of 6.7%. Moreover, on a dictation task, their model achieves a WER of 4.1%, compared to 5% for the conventional system. Similarly, the work of Prabhavalkar et al. [212] investigated a number of sequence-to-sequence methods in automatic speech recognition. These included the RNN transducer (RNN-T), attention-based models, a new model that augments the RNN-T with attention, and a Connectionist Temporal Classification (CTC) trained system that directly outputs grapheme sequences. According to their research, sequence-to-sequence approaches can compete on dictation test sets against state-of-the-art when trained on a large volume of training data. Even though deep learning has achieved state-of-the-art in speech recognition, an area that still calls for attention is speech-to-speech translation. This is because the present deep learning based speech-to-speech translation systems operate by translating sentences individually, disregarding any contextual information from preceding sentences. While research on contextual understanding has been ongoing for years, challenges persist regarding its practicality and processing efficiency, since translation typically relies on the surrounding words for context.
@@ 4.3.3 Finance
The authors in [227] proposed a Bi-LSTM-BR technique, which combined Bagging Ridge (BR) regression with Bi-LSTM as base regressors. The pre-COVID-19 and COVID-19 exchange rates of 21 currencies against the USD were predicted using the Bi-LSTM BR, and experiments showed that the proposed method outperformed ML algorithms such as DT and SVM. However, exchange rate data can be noisy and subject to non-stationarity, which can pose challenges for predictive modelling. While bagging techniques can help mitigate the effects of noise to some extent, they may struggle to capture long-term trends or sudden shifts in the data distribution, leading to suboptimal performance. To address this, Wang et al. [228] presented an approach for one-day ahead of time exchange rate prediction that concurrently considers both supervised and unsupervised deep representation features to enhance Random Subspace. Two crucial phases in the SUDF-RS technique are feature extraction and model building. First, LSTM and deep belief networks, respectively, extract the supervised and unsupervised deep representation features. To produce high-quality feature subsets, an enhanced random subspace approach was created that integrates a random forest-based feature weighting mechanism. Then, the matching base learner is trained using each feature subset, and the final outcomes are generated by averaging the outcomes of each base learner. Experiments on EUR/USD, GBP/USD and USD/JPY showed that improved accuracy was achieved using the model.
@@
In stock market prediction, several deep learning architectures have been proposed in the literature. For instance, [229], conducted a comparative study between the ANN, SVR, RF and an LSTM model. As compared to the other models discussed in the study, the LSTM model outperformed the others in predicting the closing prices of iShares MSCI United Kingdom. Similarly, using stock market historical data and financial news, Cai et al. [230] used CNN and LSTM forecasting methods to generate seven prediction models. The seven models were then combined into a single ensemble model in accordance with the ensemble learning approach to create an aggregated model. However, the accuracy of all the models’ predictions was low. Gudelek et al. [231] proposed a CNN model which used a sliding window technique and created pictures by capturing daily snapshots within the window’s bounds. With 72% accuracy, the model was able to forecast the prices for the following day and was able to generate 5 times the starting capital. In Eapen et al. [232], a CNN and Bi-LSTM model with numerous pipelines was proposed, utilising an SVM regressor model on the S&P 500 Grand Challenge dataset, and results showed enhanced prediction performance by over a factor of 6% compared to baseline models. As presented, deep learning has undeniably achieved state-of-the-art performance across various domains within finance.

### v6

@@ 4.3.1 Human Activity Recognition
Transformers: Transformers have been employed in human activity recognition to capture long-range dependencies in sensor data, enabling more accurate activity classification. In [228], a transformer-based model is enhanced by integrating a bidirectional GRU and a linear fully connected layer into each encoder block. The decoder is reduced to a fully connected layer with softmax function. This hybrid model leverages the transformer’s ability to capture long-range dependencies in sensor data and the bidirectional GRU’s strength in modeling temporal sequences, enhancing the model performance in classifying complex and collaborative activities. Similar work is reported in [229] whereby the encoder blocks of the transformer are integrated with 1D convolutional layers instead of fully connected layers, enabling the model to extract local features. Furthermore, self-supervised contrastive learning is employed to learn from unlabeled data before the model is fine-tuned using a smaller dataset. Although transformer-based models have shown remarkable performance, they suffer from high computational cost and memory requirements, making them less suitable for real-time or resource-constrained activity recognition. A study was conducted to determine the applicability of transformer models on a low-power ESP32 microcontroller [230]. The study concludes that transformer models are not suitable for tiny devices. Furthermore, the results show that a tiny transformer model with two encoder blocks achieved lower accuracy compared to the standard LSTM model, with a difference up to 14%. However, transformer models run 3x faster than LSTM, making them a viable option provided the architecture of transformers is optimized for low-power platforms.
Generative Models: A reliable deep learning model requires a large amount of training data to learn the underlying patterns of the data. However, in HAR, data collection is expensive, and the available datasets are often limited in the number of samples. To overcome this limitation, researchers employ GAN models to generate synthetic data, thus augmenting the training set. In [231], a conditional GAN is proposed to generate realistic sensor data of different activities for human activity classification. The generator consists of four 1D convolutional and maximum pooling blocks followed by two layers of LSTM and fully connected layers, while the discriminator is a standard CNN model. An enhanced conditional GAN is proposed to improve synthetic data generation for HAR [232]. This architecture integrates 1D convolutional layers with multiple fully connected networks at the generator’s input and discriminator’s output, aiming to produce higher-quality synthetic samples compared to existing CGAN models. Similar work is reported in [233], where the conditional Wasserstein GAN is employed to generate synthetic data for HAR.
In [234], sensor data is transformed into its frequency spectrum, forming an RGB-based feature space for human activity classification. Then, BiGAN is employed to generate synthetic frequency spectrum images to increase minority sample classes, thus diversifying the training set. The generator and discriminator of the proposed BiGAN are built using 2D convolutional and fully connected layers. In [235], a conditional GAN is employed to generate synthetic sensor data for therapeutic activity recognition. The generator utilizes Inception-like modules and transposed convolutional layers, while the discriminator consists of two classification pipelines: one classifies raw signals and the other pipeline classifies the Fourier-transformed signal (frequency spectrum images). Both predictions are averaged to obtain the final prediction. Table 9 lists the summary of state-of-the-art human activity recognition.
Table 9: Summary of state-of-the-art human activity recognition.
Ref. Description Results (Datasets and Metrics)
[219], 2023 This study proposed an Inception inspired model, combining convolutional layers with GRU. The convolutional blocks use multiple-sized filters to extract multiscale feature representations MHEALTH
Accuracy: 99.25%
PAMAP2
Accuracy: 97.64%
[223], 2023 This study proposed a model with three parallel feature learning pipelines, each pipeline has two sub-feature learning pipelines consisting of convolutional layers and bidirectional LSTM, and channel-wise attention before the classifier PAMAP2
Accuracy: 98.52%
WISDM
Accuracy: 97.90%
[225], 2024 This study proposed A lightweight model consists four convolutional blocks integrated with a sequential channel-spatial module and two LSTM layers PAMAP2
Accuracy: 98.52%
WISDM
Accuracy: 97.90%
[229], 2024 A modified transformer integrated with convolutional layer instead of fully connected layer UCI-HAR
Accuracy: 95.49%
Skoda
Accuracy: 87.88%
Mhealth
Accuracy: 98.43%
[233], 2024 This study employed a conditional Wasserstein GAN to generate synthetic human activity recognition accelerometry signals Local Ulster University dataset
Accuracy: 0.7453
[234], 2024 This study converted sensor data into its frequency spectrum to create an RGB-based feature space, employed BiGAN to generate synthetic frequency spectrum images to balance minority classes, and used 2D convolutional and fully connected layers in the generator and discriminator Up-Fall
Accuracy: 99.1%
Opportunity
Accuracy: 86.8%
WISDM
Accuracy: 99.12%
[235], 2025 This study employed a conditional GAN to generate synthetic sensor data for therapeutic activity recognition which has Inception-like generator and a discriminator with two classification pipelines, one for raw signals and the other for Fourier-transformed signals Local Sharif University of Technology dataset
Avg. F1-score: 0.897
@@ 4.3.2 Speech Recognition
Speech, as the primary mode of human communication, has captivated researchers for over five decades, especially since the inception of AI [236]. From the earliest endeavors to understand and replicate the complexities of human speech, to contemporary advancements leveraging cutting-edge technologies, the quest for accurate and efficient speech recognition systems has been relentless. In recent years, the emergence of deep learning techniques has revolutionized the speech recognition field. Deep learning has demonstrated unparalleled success in processing and extracting intricate patterns from vast amount of data. When applied to the realm of speech recognition, deep learning have surpassed traditional approaches by learning intricate features directly from raw audio signals, circumventing the need for handcrafted features and complex preprocessing pipelines. This paradigm shift has significantly advanced the state-of-the-art in speech recognition, enabling systems to achieve unprecedented levels of accuracy and robustness across various languages, accents, and environmental conditions. Generally, deep learning has been extended to other essential applications of speech recognition, such as speaker identification [237, 238], emotion recognition [239], language identification [240], accent recognition [241], age recognition [242] and gender recognition [243], among many others.
@@
Sequence-to-Sequence Models: In recent times, CNN and RNNs have been leveraged for automatic speech recognition in order to consider a longer or variable temporal window for context information extraction [247]. Generally, CNNs are well-suited for capturing local patterns and hierarchical features in data, making them effective for modelling acoustic features in speech. By directly learning features from raw speech signals, CNNs bypassed the need for handcrafted features used in traditional GMM-HMM systems. Additionally, CNNs can capture long-range dependencies in the data, which is crucial for understanding the context of speech. Likewise, the RNNs are suitable choice for exploring extended temporal context information in one processing level for feature extraction and modelling.
@@
Based on this, several researchers have proposed the use of both CNN and variants of RNNs for automatic speech recognition and for other speech related tasks. For instance, [248], used CNN to classify speech emotions and benchmarked on a dataset consisting of seven classes (anger, disgust, fear, happiness, neutral, sadness, and surprise). However, CNNs lack the ability to model temporal dependencies explicitly. In speech recognition, understanding the temporal context of speech is essential for accurate transcription. Furthermore, speech signals are inherently sequential, and information from previous time steps is crucial for understanding the current speech segment. CNNs, by design, do not inherently capture this sequential nature. For this reason, variants of RNNs have been leveraged to collect extended contexts in speeches. This is because RNNs are designed to model sequential data by maintaining hidden states that capture information from previous time steps. This allows them to capture temporal dependencies effectively, making them well-suited for ASR tasks. In [249], the authors evaluated the performance of RNN, LSTM, and GRU on a popular benchmark speech dataset (ED-LIUM). The results show that LSTM achieved the best word error rate, while the GRU optimization was faster and achieved word error rate close to that of LSTM.
@@
However, RNN architectures process input sequences sequentially, which limits their ability to capture global context information effectively. As a result, they may struggle to understand the entire context of a spoken utterance, leading to lower transcription accuracy, particularly in tasks requiring understanding beyond local dependencies. Furthermore, most CNN and RNN automatic speech recognition systems consist of separate acoustic, pronunciation, and language modelling components that are trained independently. Usually, the acoustic model bootstraps from an existing model that is used for alignment to train it to recognize context dependent (CD) states or phonemes. The pronunciation model, curated by expert linguists, maps the sequences of phonemes produced by the acoustic model into word sequences. For this reason, Sequence-to-Sequence models are being proposed in automatic speech recognition to train the acoustic, pronunciation, and language modelling components jointly in a single system [250]. Sequence-to-Sequence models in automatic speech recognition are a class of models that aim to directly transcribe an input sequence of acoustic features such as speech spectrograms or Mel-frequency cepstral coefficients into a sequence of characters or words representing the recognized speech. There have been various sequence-to-sequence models explored in the literature, including Recurrent Neural Network Transducer (RNN-T) [251], Listen, Attend and Spell (LAS) [252], Neural Transducer [253], Monotonic Alignments [254] and Recurrent Neural Aligner (RNA) [255].
@@
In [258], the authors explored various structural and optimization enhancements to their LAS Sequence to Sequence model, resulting in significant performance improvements. They introduce several structural enhancements, including the utilization of word piece models instead of graphemes and the incorporation of a multi-head attention architecture, which outperforms the commonly used single-head attention mechanism. Additionally, they investigate optimization techniques such as synchronous training, scheduled sampling, label smoothing, and minimum word error rate optimization, all of which demonstrate improvements in accuracy. The authors presented experimental results utilizing a unidirectional LSTM encoder for streaming recognition. On a 12,500-hour voice search task, they observe a decrease in Word Error Rate (WER) from 9.2% to 5.6% with the proposed changes, while the best-performing conventional system achieves a WER of 6.7%. Moreover, on a dictation task, their model achieves a WER of 4.1%, compared to 5% for the conventional system. Similarly, the work of Prabhavalkar et al. [250] investigated a number of sequence-to-sequence methods in automatic speech recognition. These included the RNN transducer (RNN-T), attention-based models, a new model that augments the RNN-T with attention, and a Connectionist Temporal Classification (CTC) trained system that directly outputs grapheme sequences. According to their research, sequence-to-sequence approaches can compete on dictation test sets against state-of-the-art when trained on a large volume of training data.
Transformers: Transformers have become the basis of state-of-the-art models in speech recognition. Wav2Vec 2.0 reads raw audio signals using multilayer convolutional encoder to generate latent speech representations [259]. The feature representation is subsequently fed to a transformer to capture the contextual information in the data. The transformer uses a convolutional layer for the positional encoding. Furthermore, the authors introduced vector quantization technique to convert the feature encoder output to discrete codes which are then used for contrastive learning during pretraining. Whisper is a robust speech recognition model proposed by OpenAI, trained on 680,000 hours of labeled audio data [260]. Of these, 117,000 hours cover 96 non-English languages and 125,000 hours consist of non-English to English translation data. The model is based on encoder-decoder transformer architecture, where the input data is first processed using convolutional layers followed by the positional encoding. Then, the input is fed into the encoder, which consists of a series of transformer blocks, each including multi-head self-attention mechanisms. The encoder’s output is then fed into each decode transformer block to generate the final output. The decoder’s transformer blocks consist of multi-head self-attention and cross attention mechanisms.Conformer is a Transformer-based encoder, combining self-attention with convolutional layers to better capture both global and local features in speech signals [261]. The encoder provides a balance approach leveraging the strengths of convolution and self-attention mechanism. Conformer is often used with CTC or RNN-T, or other decoding mechanisms.
Auto-regressive decoding in transformer-based models is computationally expensive, as it requires repeated processing of the complete encoded speech, resulting in slow operation. In [262], Aligner-Encoder, a transformer-based encoder is proposed that perform internal alignment between audio and text, eliminating the need for complex decoding steps and complex dynamic programming during training. The approach simplifies the model architecture using a lightweight decoder without cross attention that processes embedding frames sequentially until an end-of-message token is generated. In [263], the Spike Window Decoding (SWD) algorithm is introduced that leverages the spike property of CTC outputs where each spike represents a strong signal indicating the presence of a specific token. By focusing on these spikes within a fixed window, SWD reduces the complexity of decoding, enabling faster and more efficient processing. Although transformer-based models are highly effective, they often capture highly entangled feature representations, leading to a lack of clear interpretability. In [264], the authors proposed the Disentangled-Transformer to enhance the interpretability of transformer-based models. The transformer disentangles the internal representations into sub-embeddings based on the various temporal characteristics of the speech signals. To this end, the loss function is improved by introducing time-invariant regularization terms for each time frame. A state-of-the-art speech recognition model called Samba-ASR is proposed utilizing Mamba architecture that is based on state-space-models [265]. This approach overcomes the limitations of transformer-based models such as inability to handle long-range dependencies and their quadratic scaling with input length, by efficiently capturing both local and global temporal dependencies through state-space-dynamics. Table 10 lists the summary of state-of-the-art speech recognition applications.
Table 10: Summary of state-of-the-art speech recognition.
Ref. Description Results (Datasets and Metrics)
[260], 2022 This study proposed Whisper an automatic speech recognition (ASR) system that transcribes and translates spoken language using a robust deep learning model trained on diverse multilingual and multitask datasets Various datasets (first three datasets only)
LibriSpeech Clean
WER: 2.7
Artie
WER: 6.2
Common Voice
WER: 9.0
[262], 2025 This study introduced Aligner-Encoder, a transformer-based automatic speech recognition model that internally aligns audio-text representations before decoding, enabling a simpler and more efficient architecture trained with frame-wise cross-entropy loss LibriSpeech-960H
WER: 2.2
Voice Search
WER: 3.7
YouTube videos
WER: 7.3
[263], 2025 An algorithm that leverages spike property of CTC outputs to reduce the complexity of WFST decoding process AISHELL-1
CER: 3.89
In House
CER: 2.09
[264], 2024 A transformer-based model architecture designed to disentangle internal representations and enhance model explainability. LibriSpeech-100H
DER: 8.1
LibriMix 1
DER: 5.7
LibriMix 2
DER: 6.9
LibriMix 3
DER: 2.5
LibriMix 4
DER: 5.6
[265], 2025 A speech recognition model based on state-space-models GigaSpeech
WER: 9.12
LibriSpeech Clean
WER: 1.17
LS Other
WER: 2.48
SPGISpeech
WER: 1.84
@@ 4.3.3 Finance
The authors in [272] proposed a Bi-LSTM-BR technique, which combined Bagging Ridge (BR) regression with Bi-LSTM as base regressors. The pre-COVID-19 and COVID-19 exchange rates of 21 currencies against the USD were predicted using the Bi-LSTM BR, and experiments showed that the proposed method outperformed ML algorithms such as DT and SVM. However, exchange rate data can be noisy and subject to non-stationarity, which can pose challenges for predictive modelling. While bagging techniques can help mitigate the effects of noise to some extent, they may struggle to capture long-term trends or sudden shifts in the data distribution, leading to suboptimal performance. To address this, Wang et al. [273] presented an approach for one-day ahead of time exchange rate prediction that concurrently considers both supervised and unsupervised deep representation features to enhance Random Subspace. Two crucial phases in the SUDF-RS technique are feature extraction and model building. First, LSTM and deep belief networks, respectively, extract the supervised and unsupervised deep representation features. To produce high-quality feature subsets, an enhanced random subspace approach was created that integrates a random forest-based feature weighting mechanism. Then, the matching base learner is trained using each feature subset, and the final outputs are generated by averaging the outcomes of each base learner. Experiments on EUR/USD, GBP/USD and USD/JPY showed that improved accuracy was achieved using the model.
@@
In stock market prediction, several deep learning architectures have been proposed in the literature. For instance, [274], conducted a comparative study between the ANN, SVR, RF and an LSTM model. As compared to the other models discussed in the study, the LSTM model outperformed the others in predicting the closing prices of iShares MSCI United Kingdom. Similarly, using stock market historical data and financial news, Cai et al. [275] used CNN and LSTM forecasting methods to generate seven prediction models. The seven models were then combined into a single ensemble model in accordance with the ensemble learning approach to create an aggregated model. However, the accuracy of all the models’ predictions was low. Gudelek et al. [276] proposed a CNN model which used a sliding window technique and created pictures by capturing daily snapshots within the window’s bounds. With 72% accuracy, the model was able to forecast the prices for the following day and was able to generate 5 times the starting capital. In Eapen et al. [277], a CNN and Bi-LSTM model with numerous pipelines was proposed, utilising an SVM regressor model on the S&P 500 Grand Challenge dataset, and results showed enhanced prediction performance by over a factor of 6% compared to baseline models. As presented, deep learning has undeniably achieved state-of-the-art performance across various domains within finance. In [278], a multi-modal deep learning model is proposed for stock market trend prediction by integrating daily stock prices, technical indicators and sentiment from daily news headlines. The model architecture consists of BERT-based branch fine-tuned on financial news to capture textual sentiment and an LSTM branch that captures temporal patterns in the time series data, including stock prices and technical indicators. Both feature representations are combined through concatenation and then passed to a fully connected layer for predictions.
A study was conducted to analyze the performance of Prophet, LSTM and Transformer. Prophet is a parametric, additive regression model based on time series decomposition [279]. The LSTM model consists of two hidden layers, with a hidden size of 64 and a linear fully connected layer is used to generate the output. The Transformer model follows the original transformer architecture. The results show that transformer outperformed LSTM and Prophet in terms of forecasting accuracy, particularly for datasets with complex temporal dependencies. However, transformer models have high memory usage and quadratic complexity due to self-attention mechanism, which makes them inefficient for very long sequence time-series forecasting. To overcome the limitations, a StockFormer is proposed which is based on Informer [280]. Informer is a transformer that was designed for time-series forecasting, utilizing the ProbSparse self-attention mechanism which attends to only the most important queries for each key, reducing the computational complexity [281]. Furthermore, self-attention distilling technique is introduced to reduce redundancy and highlighting crucial information in the feature maps, helping the model to focus more on the relevant parts of the input sequence. In [282], a transformer-based model is proposed for stock trend prediction using limit order book data (LOB). The model consists of a series of transformer LOB (TLOB) blocks which leverage dual attention mechanisms to capture both spatial and temporal dependencies, allowing the model to adaptively focus on the market microstructure of the input data. Furthermore, a multilayer of perceptron LOB (MPLOB) is introduced, comprising two blocks of two fully connected layers which operate along the feature and temporal dimensions. Table 11 lists the summary of state-of-the-art finance applications.
Table 11: Summary of state-of-the-art finance applications.
Ref. Description Results (Datasets and Metrics)
[273], 2023 This study introduces SUDF-RS technique for one-day-ahead exchange rate prediction by combining supervised and unsupervised deep representation features with an enhanced random subspace approach, leveraging LSTM and deep belief networks for feature extraction and using a random forest-based weighting mechanism to improve prediction accuracy Data collected from investing.com
RMSE: 0.448 (EUR/USD), 0.676 (GBP/USD), 0.374 (USD/JPY)
[278], 2024 A multi-modal deep learning model integrates stock prices, technical indicators, and news sentiment using a BERT-based branch for textual analysis and an LSTM branch for temporal patterns, combining their features for stock market trend prediction Data collected using Yahoo Finance and EODHD API
F1-Score: 0.45 (ATVI), 0.51 (AAPL), 0.48 (AMT), 0.53 (PLD), 0.50 (NDAQ), 0.45 (SCHW), 0.43 (BIO), 0.47 (JNJ)
[279], 2024 This study analyzes stock closing price prediction using LSTM, Prophet, and Transformer models American Airlines Group Inc.
MSE: 0.0085 (Transformer), 0.1972 (LSTM), 16.9321 (Prophet)
Atlantic American-Life Insurance
MSE: 0.0118 (Transformer), 0.0095 (LSTM), 17.8601 (Prophet)
[280], 2025 The study proposed a model based on Informer which utilizes ProbSparse self-attention and self-attention distilling to reduce the computation complexity Combined data from Yahoo, alphavantage.co, alpaca.markets and polygon.io
Percent profit: 1.7550
[282], 2025 The study proposed a transformer-based model with dual attention mechanisms and multilayer of perceptrons for feature mixing along feature and temporal dimensions FI-2010
F1-score: 92.81
TSLA
F1-score: 60.50
INTC
F1-score: 80.15